# Contactless Fingerprint Recognition Using 3D Graph Matching

### v1

@@ III-B3 Training Data
III-C1 3D Minutiae Feature
The direct outputs of the feature extraction network are 2D. They need to be transformed into 3D space to fit for the 3D graph matching. Fig. 5 shows some examples of 3D minutiae mapped in 3D space. Compared with 2D minutiae, 3D minutiae have an extra dimension, which needs to be calculated out.
The 3D depth $z$ is recovered from the gradient results $g_{x}$ and $g_{y}$ using numerical integration, as the gradient is the differential of the surface depth. Detailed calculation procedures can be found in [18]. The calculated finger depth represents the $z$ coordinate, and is combined with the original image pixel $(x,y)$ to get a 3D coordinate $\mathbf{p}=(x,y,z)$.
Next, the minutiae orientation also needs to be recalculated in 3D space along with the tangential direction on the surface. As shown in Fig. 6, Assuming the 3D orientation is $(\sin\phi\cos\theta,\sin\phi\sin\theta,\cos\phi)$, where $(\theta,\phi)$ is the standard spherical coordinate. The 2D minutiae orientation can be described as $(\cos\theta,\sin\theta)$, as it is the projection of 3D orientation onto the $x-y$ plane.
On the other side, the 3D minutiae orientation should be perpendicular to the surface normal $\vec{n}$. We now have its gradient $(g_{x},g_{y})=(\partial z/\partial x,\partial z/\partial y)$ for a surface point. Based on differential geometry knowledge, we know $\vec{n}$ should be $(g_{x},g_{y},1)$. Therefore, we have:
$(\sin\phi\cos\theta,\sin\phi\sin\theta,\cos\phi)^{T}(g_{x},g_{y},1)=0.$ (6)
The 2D orientation $(\cos\theta,\sin\theta)$ and gradient $(g_{x},g_{y})$ are known from the 2D orientation and gradient outputs. Solving equation 6 can get the 3D orientation $(\theta,\phi)$.
However, directly using $(\sin\phi\cos\theta,\sin\phi\sin\theta,\cos\phi)$ to represent the orientation of the minutiae would result in a scale difference compared to the spatial coordinates $\mathbf{p}=(x,y,z)$. Hence, we need to scale the length of the minutiae orientation. The final representation of the 3D minutiae orientation is:
$\mathbf{o}=(d_{x},d_{y},d_{z})=\alpha(\sin\phi\cos\theta,\sin\phi\sin\theta,\cos\phi).$ (7)
By analyzing 3D minutiae data, we chose $\alpha=25$ empirically to balance the scale parameter. So the final representation of each 3D minutiae would be:
$\boldsymbol{m}=(\mathbf{p},\mathbf{o})=(x,y,z,d_{x},d_{y},d_{z}).$ (8)
@@
Figure 7: The architecture of our 3D graph matching network. The specific architecture of EdgeConv can be found in [45].
III-C2 Graph Matching Network
@@
Our graph matching network is inspired by [23]. We use its Topological Relation Reasoning Module (TRM) module, and simplify this module just like the classification model of DGCNN [45]. We encode each fingerprint’s 3D minutiae using DGCNN. Through EdgeConv, the network adopts a convolution-like approach, merging the 3D minutiae around each minutia in the set. In this way, a fixed-length encoding similar to the Minutia Cylinder Code (MCC) [46] can be obtained for each fingerprint. Since this method is trained on a large amount of fingerprint data, it is more robust to noise compared to MCC.
To address the over-smoothing issue mentioned in [23], we do not adopt the Feed-Forward Module (FFM) mentioned in the article, which will reduce the accuracy of our validation set. We only add residuals to the first and third EdgeConv, thereby somewhat mitigating the over-smoothing issue.
@@
The detailed architecture of our network is shown in Fig. 7. Our network consists of two parts: the spatial transformation network and the graph embedding network. The spatial transformation network predicts and corrects the posture of 3D minutiae in 3D space. The graph embedding network extracts the 3D minutiae with corrected posture into 256-dimensional fixed-length features for matching. The two network structures are similar, except that the spatial transformation network contains five layers of MLP while the graph embedding network contains only two layers.
Each part of our network consists of 5 layers of EdgeConv. The first layer maps the number of channels from 6 to 64, and the number of channels doubles in each layer. We introduce residual connections in the first and third EdgeConv layers. Afterward, the outputs of all EdgeConv layers are concatenated along the channels, resulting in an output with 1024 channels. This output undergoes maximum and average pooling for all 3D minutiae of each fingerprint, and the results are concatenated.
@@
For the spatial transformation network, since the frequency of contactless fingerprints has been unified during preprocessing, we only need to correct the translation and rotation of 3D minutiae. After passing through five MLP layers, we obtain a 6-dimensional feature. The first three dimensions correspond to the translation in three-dimensional space $\mathbf{t}=(x_{t},y_{t},z_{t})$, and the last three dimensions correspond to the rotation in the form of Euler angles in the $xyz$ order in three-dimensional space $(\theta_{t},\psi_{t},\phi_{t})$. The six dimensions are then used to calculate the three-dimensional rotation matrix $\mathbf{R}$ using Equation 9.
[…]
$\mathbf{R}_{z}=$ $\begin{bmatrix}1&0&0\\ 0&\cos{\phi_{t}}&-\sin{\phi_{t}}\\ 0&\sin{\phi_{t}}&\cos{\phi_{t}}\end{bmatrix}\mathbf{R}=\mathbf{R}_{z}\mathbf{R}_{y}\mathbf{R}_{x}$
$\mathbf{R}_{x}$, $\mathbf{R}_{y}$, and $\mathbf{R}_{z}$ represent the rotation of 3D minutiae along the $x$, $y$, and $z$ axes, respectively.
@@
Then, $\mathbf{M}_{t}$ is sent to the graph embedding network. After passing through two MLP layers, we obtain a 256-dimensional feature for the graph embedding network. Then, we send that fixed length feature to calculate loss. The loss function is the same as the loss function in [23], which is a triplet loss [47]:
@@
III-C3 Training Data
@@
In stage I, we collect a vast number of plain 2D fingerprints from NIST301 [48], NIST302 [49], CASIA-Fingerprint v5 [50], PrinsGAN [51], and SpoofGAN [52]. Then, we transfer these contact-based 2D fingerprints to 3D space to create synthetic contactless 3D fingerprints. However, given a contact-based 2D fingerprint, there are barely any methods of transferring a contact-based 2D fingerprint to a 3D fingerprint. Therefore, our research adopts a rough but efficient approximation method to project 2D fingerprint to 3D space. As Fig. 8 shows, we generally assume that the finger skin that touches the sensor is a sphere, and the contact-based 2D fingerprint is a projection of the 3D skin surface on a sphere.
@@
In stage II, the data used for finetuning is the 3D minutiae from contactless fingerprints. We use Verfinger 13.0 [44] and our feature extraction network to extract minutiae from the last 2,988 fingerprints in the UWA database [41]. Due to Verifinger failing to extract features on some fingerprints, we obtain 2,879 images. Subsequently, we manually filtered out fingerprints with low-quality depth maps, resulting in 2,779 fingerprint images for training. In these 2,779 images, each fingerprint consists of at most 6 images.
@@
Due to the higher frequency of contactless fingerprints with more minor pose variations in real-life scenarios, we sample the fingerprint pairs in the same direction three times. The fingerprint pairs of front and left poses, as well as front and right poses, are sampled twice, making the network easier to train. Therefore, we ultimately obtain 12,830 pairs of fingerprints in the UWA database for training.
@@
For the CFPose database, we adopt the last 200 fingerprints for training. We perform three-times samplings for fingerprint pairs with a yaw angle difference $\Delta{\theta}\leq 10^{\circ}$ along the y-axis, six-times samplings for those with $10^{\circ}<\Delta{\theta}\leq 45^{\circ}$, and three-times samplings for the rest, achieving data balance within the CFPose dataset as well as the entire training set. The yaw angle is estimated by the method in [11]. We ultimately obtain 4,230 pairs of fingerprints in the UWA database for training.
@@
We perform data augmentation on the input 3D minutiae in finetuning, including random 3D rotation ($x$ and $z$ axes range from $-30^{\circ}$ to $30^{\circ}$, $y$ axes range from $-90^{\circ}$ to $90^{\circ}$), random translation (all three axes range from $-300$ to $300$ pixels), and random dropping of $1/4$ of all the 3D minutiae for the fingerprints with $\Delta{\theta}\leq 10^{\circ}$. For the remaining fingerprints, we perform random translation (all three axes range from $-100$ to $100$ pixels), and random dropping of $1/6$ of all the 3D minutiae.
However, fingerprints with larger pose variations have a smaller overlap area, which misleads the network training. As a result, although the performance on EER is better, the performance on Rank-1 accuracy is worse. Thus, we only extract the fingerprints of the same orientation (1,494 pairs) to train an individual network, and fuse the matching scores of the network output obtained by this training method with the matching scores of the network output obtained by the above training method. The fusion can slightly increase the EER while achieving a higher Rank-1 accuracy than the experimental results obtained by the two networks.
If the matching score of the former is $s_{1}$, the matching score of the latter is $s_{2}$, and the fused score is $s$, then our score fusion is:
$s=\begin{cases}0.6s_{1}+0.4s_{2}&s_{2}>=0.7\\ s2&s_{2}<0.7\end{cases}.$ (14)
If we only calculates $s=0.6s_{1}+0.4s_{2}$, we can achieve a high Rank-1 accuracy, but the EER will also be high. Replacing the lower scores with $s_{2}$ results in a significant reduction of the EER.
When our method is combined with the matching scores of existing methods [11][18], our approach achieves state-of-the-art performance on the testing set in terms of EER and Rank-1 accuracy. The matching results obtained by both methods are Verifinger’s matching scores. If the Verifinger matching score is $s_{3}$, the fused score is calculated as $s_{3}/1000+s/300$.
III-D Implementation Details
III-D1 Feature Extraction
Our feature extraction network is trained on a combination of 2,988 fingerprints from the UWA dataset and 200 fingerprints from the CFPose dataset. After preprocessing the original images and computing orientation fields, periodic maps, surface depth gradients, and minutiae, we extract two 512$\times$512 patches from each original image, mask, and label. In total, there are 5,872 patches, with 5,506 from the UWA dataset and 366 from CFPose.
Since the network’s output feature sizes are 1/8 of the original image, we downsample each label to 64$\times$64 during input. During testing, we upsample the predicted results by 8 times and compute various losses against the ground truth labels.
The network is initially trained on the UWA dataset and then finetuned with the addition of the CFPose dataset. We use the Adam optimizer with a learning rate of 0.001, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.999, and a decay rate of 0. The batch size is set to 16, and the training loss is a weighted sum of various components: 1 $\times$ direction field, 60 $\times$ periodic map, 300 $\times$ gradient, 300 $\times$ minutiae score, 15 $\times$ minutiae horizontal position, 15 $\times$ minutiae vertical position, 15 $\times$ minutiae orientation.
III-D2 3D Graph Matching
Our 3D graph matching network is trained using the Adam optimizer with optimizer parameters: learning rate$=0.001$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay$=5e-4$.
We first train only a graph embedding network on NIST301 [48], NIST302 [49], CASIA-Fingerprint v5 [50], PrinsGAN [51], and SpoofGAN [52] because our contact-based fingerprints projected into 3D space do not require pose correction. The training pairs in the UWA database for finetune are also in the pretraining pairs. The total training pairs are 998,175. There are 2,396 fingerprints in NIST302 [49] used for validation. The details for finetune are shown in III-C3.
For batch training, we pad the number of 3D minutiae to 200 on each fingerprint with zeros during pretraining, and 400 during finetuning. The batch size of pretraining is 128 and 64 for finetuning. We train 80 epochs for pretraining and 100 epochs each network for finetuning.
Our experiments are conducted on an Nvidia 4090 GPU and an Intel Xeon Gold 6133 CPU @ 2.50GHz.
IV Experiment Results
IV-A Datasets
TABLE II: Databases used in this paper
Database
IDs
Contact-based
/Contactless
Fingerprints
Finger
Pose
Training
Data
Evaluation
Experiment
Evaluation
Data
NIST301[48] 469 3,053/0 -
All for pretraining
Matching Network
NIST302[49] 2,000 23,155/0 -
All for pretraining
Matching Network
SpoofGAN*[52] 1,500 31,500/0 -
All for pretraining
Matching Network
PrintsGAN*[51] 34,992 524,880/0 -
75000 for pretraining
Matching Network
CASIA
FingerprintV5[50]
500 20,000/0 -
All for pretraining
Matching Network
UWA[41] 1,500 0/8,958 Left, Front, Right
The last 2,988 fingerprints for
training Feature Network and
finetuning Matching Network
Feature Network
Matching Network
The last 2,988 fingerprints for
Feature Network evaluation
& The first 5,970 fingerprints
for Matching experiments
CFPose[11] 140 0/1,400 Random 10 Poses
The last 200 fingerprints for
finetuning Matching Network
Feature Network
Matching Network
The last 200 fingerprint for
Feature Network evaluation
& The first 1,200 fingerprints
for Matching experiments
  • *: Synthetic data.
As listed in Table II, seven datasets are used in the experiments: the primary datasets we use are the UWA Benchmark 3D/2D Fingerprint Database [41] and the CFPose Fingerprint Database [11]. NIST301 [48], NIST302 [49], CASIA-Fingerprint v5 [50], PrinsGAN [51], and SpoofGAN [52] are used for pretraining the matching network but not for evaluation.
UWA Database [41] contains 8,958 contactless fingerprints with pose variations. The ground-truth depth maps are reconstructed using the shape-from-silhouette algorithm in [18]. The ground-truth orientation, period, and minutiae are obtained by VeriFinger [44]. The last 2,988 fingerprints are used to train and test the network, and the first 5,970 fingerprints are used for matching experiments.
CFPose [11] contains 1,400 contactless fingerprints with random pose variations. The ground-truth depth maps are reconstructed using the ellipsoid model algorithm in [11]. The last 200 fingerprints are used to train and test the network, and the first 1,200 are used for matching experiments.
IV-B Evaluation of Feature Network
The proposed feature extraction network is evaluated on the testing set from UWA and CFPose databases. The network has four outputs: gradient, orientation, ridge, and minutiae, and they are all evaluated. For the UWA dataset, evaluations are conducted on the front and side poses. For the CFPose dataset, each finger has 10 random poses, and we categorize those fingerprints with $\theta>30^{\circ}$ as side poses.
Table III shows the evaluation results of orientation, ridge period, gradient, and reconstructed depth. The depth is reconstructed using the estimated gradient by the reconstruction method in [18]. The units of errors are: orientation in degree, period and depth in pixel, gradient in pixel per pixel. If converted to real-life size, a pixel would equal to 0.0508mm. So the reconstruction error is approximately 2mm in real-life, which is very tiny compared to the size of a finger.
TABLE III: Evaluation results of the feature extraction network on the UWA and CFPose database
Database Poses
Orientation
Period
Gradient
(Pixel/Pixel)
Depth

### v2

@@ 3.2 Fingerprint Feature Extraction
Fig. 4 illustrates the calculation process of the 3D minutia, where $\phi=\frac{\pi}{2}-\varphi$. To make the scales of minutiae direction and position comparable, we multiply the minutiae direction by a constant $n$, where $n$ is set to 25.
@@
Figure 5: The architecture of our 3D graph matching network. The specific architecture of EdgeConv can be found in [wang2019dynamic].
@@
Graph Matching Network. Our graph matching network is inspired by [su2023gnn]. We use its Topological Relation Reasoning Module (TRM) module, and simplify this module just like the classification model of DGCNN [wang2019dynamic]. Through EdgeConv, the network adopts a convolution-like approach, merging the 3D minutiae around each minutia in the set.
@@
The detailed architecture of our network is shown in Fig. 5. Our network consists of two parts: the spatial transformation network and the graph embedding network. The spatial transformation network predicts and corrects the posture of 3D minutiae in 3D space. The graph embedding network extracts the 3D minutiae with corrected posture and deep features into 256-dimensional fixed-length features for matching. We project the 3D minutiae to 128-dim and sum with the deep features into a 128-dimensional input for the graph embedding network.
@@
We adopt a spatial transformation network, since the frequency of contactless fingerprints has been unified during preprocessing. We only need to correct the translation and rotation of 3D minutiae. After passing through five MLP layers, we obtain a 6-dimensional feature. The first three dimensions correspond to the translation in three-dimensional space $\mathbf{t}=(x_{t},y_{t},z_{t})$, and the last three dimensions correspond to the rotation in the form of Euler angles in the xyzxyz order in three-dimensional space $(\theta_{t},\psi_{t},\phi_{t})$. The six dimensions are then used to calculate the three-dimensional rotation matrix $\mathbf{R}$ using $\mathbf{R}=\mathbf{R}_{z}^{\phi_{t}}\mathbf{R}_{y}^{\psi_{t}}\mathbf{R}_{x}^{\theta_{t}}$, in which $\mathbf{R}_{x}^{\theta_{t}}$, $\mathbf{R}_{y}^{\psi_{t}}$, and $\mathbf{R}_{z}^{\phi_{t}}$ represent the rotation of 3D minutiae along the $x$, $y$, and $z$ axes, respectively.
@@
Then, $\mathbf{M}_{t}$ and 3D feature $\mathbf{F}$ are fused to 128-dimensional, and are sent to the graph embedding network. After passing through two MLP layers, we obtain a 256-dimensional feature for the graph embedding network. Then, we send fixed length feature to calculate loss. The loss function is the same as the loss function in [su2023gnn], which is a triplet loss [schroff2015facenet]:
@@
Similar to the method of Cui et al. [cui2023monocular], we use our depth prediction results to unwarp the fingerprint images and perform fingerprint matching using VeriFinger. We then fuse the matching score with the graph matching similarity score to achieve state-of-the-art matching performance. If the VeriFinger matching score is denoted as $s_{1}$ and the graph matching score as $s_{2}$, the fused score is calculated as $s_{1}/1000+s_{2}/40$.
4 Experiment Results
4.1 Datasets
The primary contactless datasets we use are UWA Benchmark 3D/2D Fingerprint Database [zhou2014benchmark], PolyU 3D+ Database [lin2017tetrahedron], CFPose Fingerprint Database [tan2020towards] and ZJU Databases [grosz2022contact].
UWA Database [zhou2014benchmark] contains 8,958 contactless fingerprints with pose variations. The ground-truth depth maps are reconstructed using the shape-from-silhouette algorithm in [cui2023monocular]. The ground-truth orientation, period are obtained by VeriFinger 13.1 [VeriFinger]. The last 2,988 fingerprints are used to train and validate the 3D feature extraction and graph embedding network, and the first 5,970 fingerprints are used for feature extraction and matching experiments. Since our 3D feature network needs to extract fingerprint minutiae, and the minutiae predicted by VeriFinger are often not accurate enough, we randomly selected 276 images from the training set and manually adjusted the minutiae predicted by VeriFinger as ground truth. These adjusted minutiae are used to fine-tune the minutiae prediction part of the 3D feature network.
PolyU 3D+ Database [lin2017tetrahedron] We utilized the 2,016 ground truth depth maps and corresponding images from the first session for depth estimation experiments. This dataset, distinct from the UWA database, provides ground truth depth, better illustrating the effectiveness of our method.
CFPose [tan2020towards] contains 1,400 contactless fingerprints with random pose variations. The ground-truth depth maps are reconstructed using the ellipsoid model algorithm in [tan2020towards]. The last 200 fingerprints are used to train and validate the 3D feature extraction and graph embedding network, and the first 1,200 are used for matching experiments. The 40 ground truth minutiae manually annotated by Tan and Kumar [tan2020towards] are also used to fine-tune the minutiae component of the 3D feature network.
ZJU Database [grosz2022contact] contains 9,888 contactless fingerprints. We manually annotated the first 912 images for training the segmentation network.
Other Database 39,461 ids and 602,577 images in NIST301 [sd301], NIST302 [sd302], CASIA-Fingerprint v5 [casiav5], PrinsGAN [engelsma2022printsgan], and SpoofGAN [grosz2022spoofgan] are also used for pretraining the matching network but not for evaluation.
4.2 Implementation details
Our feature extraction network is trained on 2,988 fingerprints from the UWA dataset. After preprocessing the original images and computing orientation fields, periodic maps, surface depth gradients, and minutiae, we extract two 512×512 patches from each original image, mask, and label. In total, there are 24,014 patches. The network is initially trained on the UWA dataset with minutiae extracted by VeriFinger and then finetuned with the addition 2,670 patches of the UWA and CFPose dataset with ground truth minutiae only for the Minutiae Decoder and Local MinNet. We use the Adam optimizer with a learning rate of 0.0001, $\beta_{1}$ of 0.9, $\beta_{2}$ of 0.999, and a decay rate of 0. To locate minutiae, we set $\tau_{init}=0.4$ and $\tau_{area}=45$ for minutiae extraction experiment and $\tau_{init}=0.2$ and $\tau_{area}=30$ for matching.
Our 3D graph matching network is trained using the Adam optimizer with optimizer parameters: learning rate= 0.001, $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay= 5e-4. For batch training, we pad the number of 3D feature to 200 on each fingerprint with zeros during pretraining and finetuning. The batch size is 128. We train 80 epochs for pretraining and 100 epochs each network for finetuning. During pre-training, we project 2D fingerprint minutiae onto a 3D sphere to obtain a 6-dimensional input and use DMD [pan2024latent] to get 2D fixed length deep feature around minutiae.
Our experiments are conducted on two Nvidia 4090 GPUs and an Intel Xeon Gold 6133 CPU @ 2.50GHz.
4.3 Evaluation of Feature Network
The proposed feature extraction network is evaluated on the testing set from UWA and PolyU 3D+ databases for depth estimation and CFPose database for minutiae extraction. We do not evaluated orientation and period because the orientation and period are extracted by VeriFinger and may not be accurate. The depth in the UWA database, while not ground truth, is more accurate than monocular estimation due to multi-view reconstruction. Additionally, datasets with ground truth depth are relatively scarce. For the depth estimation, evaluations are conducted from gradient accuracy and depth accuracy. For minutiae extraction, we follow the evaluation protocol in [tan2020towards] for testing.
Table 1 shows the error of gradient and reconstructed depth. The depth is reconstructed using the estimated gradient by the reconstruction method in [cui2023monocular]. The units of errors are: gradient in mm per pixel, depth in mm.
Figure 6: Examples of depth estimated by Cui et al.’s method [cui2023monocular] and ours.
Database Method
Gradient
(mm/Pixel)
Depth
($m$)
UWA
Cui et al. [cui2023monocular]
0.0321 1.4629
Proposed
0.0274 1.1154
PolyU 3D+
Cui et al. [cui2023monocular]
0.0290 2.9577
Proposed
0.0208 1.3455
Table 1: Evaluation results of the depth network on the UWA and CFPose database
Fig. 6 provides an example of comparing the results of depth prediction between our method and Cui et al.’s method. Although Cui et al.’s method can predict relatively smooth depth on UWA, it struggles to capture depth variations. For instance, in the UWA example image, the depth changes sharply and shows a decreasing trend at the bottom, but Cui et al.’s method still predicts an increasing depth trend. Additionally, the PolyU 3D+ example demonstrates that Cui et al.’s method fails to generalize to data outside the training set, resulting in failed predictions.
Database Method Precision Recall F1-score
CFPose
VeriFinger
63.8% 79.1% 70.1%
Tan and Kumar [tan2020towards]
79.0% 76.4% 77.2%
Proposed
80.6% 74.1% 76.7%
UWA
VeriFinger
51.7% 88.1% 64.7%
Tan and Kumar [tan2020towards]
69.1% 74.1% 71.0%
Proposed
74.4% 72.5% 73.0%
  • $\star$: [tan2020towards] is intentionally designed for minutiae extraction, while our’s is a multi-feature network.
Table 2: Evaluation results of minutiae extraction on the UWA and CFPose database
We freeze the other parameters and only use the annotated ground truth minutiae to train the Minutiae Decoder and Local MinNet, and tested our minutiae prediction results on the ground truth dataset provided by Tan and Kumar [tan2020towards]. A minutiae prediction is considered correct if its position deviated from the ground truth by 16 pixels. As Table. 2 shown, the experimental results show that our method achieves a significantly higher accuracy in minutiae prediction compared to VeriFinger, demonstrating that our approach can reliably extract contactless fingerprint minutiae. By incorporating more training data from the UWA dataset, our method enables a single multi-task network to achieve higher minutiae prediction accuracy on the UWA dataset compared to ContactlessNet proposed by Tan and Kumar [tan2020towards].
Evaluation of these features is just a references here to exam whether the network has produced reasonable features for the matching algorithm.
4.4 Evaluation of Fingerprint Matching
Matching experiments are performed on UWA [zhou2014benchmark] and CFPose [tan2020towards]. Contactless fingerprint matchings are performed, as the proposed method aims at utilizing 3D features to improve contactless fingerprint matching performances, which is in contrast to many previous methods that aim at contactless-contact matching. These CL2C methods also do not solve the pose variation issue in CL2CL matching. We compute similarity of two fingerprints through the product of corresponding embedding, which is
$s(F_{i},F_{j})=<M_{F_{i}},M_{F_{j}}>$ (6)
where $F_{i}$ and $F_{j}$ denote two fingerprints, $M_{F_{i}}$ and $M_{F_{j}}$ are the fingerprint feature embedding extracted by our graph matching network.
We compare our matching results with the current SOTA contactless-concactless fingerprint matching algorithms, which are shown in Table 3. We conduct experiments on UWA according to Cui et al.’s [cui2023monocular] protocol and on CFPose and ZJU according to Dong and Kumar’s [dong2023syn] protocol. The DET curve on UWA database is shown in Fig. 7.
Figure 7: DET curves on UWA database
We compare our method with classical fingerprint recognition approaches [VeriFinger][rohwedder2023benchmarking] and the recent deep learning methods [cui2023monocular][tan2020towards]. Using the same pretrained and finetuning data, we train on an open-source DeepPrint [rohwedder2023benchmarking] and perform matching experiments. We reproduce [cui2023monocular][tan2020towards] for comparison. Due to the method [cui2023monocular] failing to predict reliable depth on the ZJU and CFPose databases, and also no pose variation for method [tan2020towards] to correct on ZJU, we do not reproduce them.
Method EER
CFPose[tan2020towards] UWA[zhou2014benchmark] ZJU[grosz2022contact]
DeepPrint [rohwedder2023benchmarking] 5.72% 28.20% 6.43%
VeriFinger [VeriFinger] 5.44% 20.32% 0.53%
Tan and Kumar [tan2020towards] 1.79% 15.22% -
Cui et al. [cui2023monocular] - 13.40% -
Dong and Kumar [dong2023syn] 1.12% 13.84% 0.48%
Proposed 0.91% 11.16% 0.44%
Proposed 0.89% 9.23% 0.37%
  • *: A different test protocol with ours.
  • $\dagger$}: Only unwarp by our predicted depth.
Table 3: Matching results on three databases
First, following the approach of Cui et al. [cui2023monocular], we unwarp the fingerprints based on the depth predicted by our 3D feature net and perform matching using VeriFinger 13.1. At this stage, we have already achieved state-of-the-art EER on all three datasets. To demonstrate the integration of 3D features into fingerprint matching, we fuse our graph matching scores with those from VeriFinger, resulting in a significant boost in matching performance. Notably, when tested under the protocol of another CL2CL method [dong2023syn], our EER is 12.07% compare to there 13.84%, further proving our method’s effectiveness. Our method better identifies fingerprints with large pose variations especially on the UWA dataset with significant pose variations. This improvement shows our method effectively models large pose changes in contactless fingerprints, adding more matching information.
Fig. 8 shows genuine and impostor examples of our matching scores compared to VeriFinger’s. Our graph matching method can better capture the internal information between genuine contactless fingerprints with exaggerated poses. So, it can get a high score on genuine pairs and low score on imposter pairs with similar textures, while VeriFinger is more likely to be misled by fake matching fingerprints with similar textures.
Figure 8: Two pairs of matching result. (a) Genuine matching fingerprints have viewpoint difference. (b) Impostor matching fingerprints have viewpoint difference.
5 Conclusion
In this paper, we propose a novel contactless fingerprint recognition method that uses 3D graph matching to implement a 3D matching algorithm for contactless fingerprints. In contrast, previous methods only match fingerprints in 2D space and neglect the 3D information during matching. Experiment results on three databases prove that our method reaches state-of-the-art matching performances, and can successfully match contactless fingerprints across different poses.
Supplementary Material
Appendix A Implementation Details
A.1 Preprocessing
Contrast. We initially selected the first phalanx as foreground region. The maximum convex hull of these regions is identified as a mask, followed by the application of contrast-limited adaptive histogram (CLAHE) [zuiderveld1994contrast] to normalize the contrast of the segmented images while minimizing the effect of noise. The grid size of CLAHE is empirically set to 60 pixels.
Pose. The finger’s shape allows us to directly obtain the centerline of the finger through the contour in the preprocessing stage. Rotating the finger’s centerline to a vertical position eliminates rotations along the $z$-axis, providing a certain level of normalization.
Scale. The scale of contactless fingerprint capture varies significantly, leading to substantial variations in the frequency of captured fingerprint ridge lines. It is necessary to normalize ridge frequency, which could affect fingerprint recognition. Thus, the ridge frequency in the central region of the fingerprint is balanced to 10 pixels across fingerprint images, avoiding affecting the subsequential feature extraction and matching steps. The average ridge period within the central area is computed using a traditional algorithm [hong1998fingerprint]. However, due to the low quality of contactless fingerprints, the extracted period can sometimes be too large or too small. When the scaling factor calculated based on the frequency is greater than 1.3 or less than 1, we consider the frequency extraction to have failed. In this case, we compute the scaling factor based on the mask, setting it to three-quarters of the fingerprint image width.
The final output fingerprint size is $640\times 480$.
A.2 Graph Embedding Network
@@
In stage I, we collect a vast number of plain 2D fingerprints from NIST301 [sd301], NIST302 [sd302], CASIA-Fingerprint v5 [casiav5], PrinsGAN [engelsma2022printsgan], and SpoofGAN [grosz2022spoofgan]. Then, we transfer these contact-based 2D fingerprints to 3D space to create synthetic contactless 3D fingerprints. However, given a contact-based 2D fingerprint, there are barely any methods of transferring a contact-based 2D fingerprint to a 3D fingerprint. Therefore, our research adopts a rough but efficient approximation method to project 2D fingerprint to 3D space. As Fig. 9 shows, we generally assume that the finger skin that touches the sensor is a sphere, and the contact-based 2D fingerprint is a projection of the 3D skin surface on a sphere.
Figure 10: DET curves on CFPose (left) and ZJU (right) database
@@
In stage II, the data used for finetuning is the 3D minutiae based feature points from contactless fingerprints. We extract some minutiae based feature points via our 3D feature net. Specifically, we adjust the recall threshold $\tau_{init}$ for minutiae. $\tau_{init}$ is 0.4 when recalling minutiae and 0.2 when recalling 3D minutiae based feature points.
@@
Due to the higher frequency of contactless fingerprints with more minor pose variations in real-life scenarios, we sample the fingerprint pairs in the same direction three times. The fingerprint pairs of front and left poses, as well as front and right poses, are sampled twice, making the network easier to train.
@@
For the CFPose database, we adopt the last 200 fingerprints for training. We perform three-times samplings for fingerprint pairs with a yaw angle difference $\Delta{\theta}\leq 10^{\circ}$ along the y-axis, six-times samplings for those with $10^{\circ}<\Delta{\theta}\leq 45^{\circ}$, and three-times samplings for the rest, achieving data balance within the CFPose dataset as well as the entire training set. The yaw angle is estimated by the method in [tan2020towards].
@@
We perform data augmentation on the input 3D minutiae in finetuning, including random 3D rotation (all axes range from $15^{\circ}$ to $15^{\circ}$) the fingerprints with $\Delta{\theta}\leq 10^{\circ}$. For the all fingerprints, we also perform random translation (all three axes range from 100-100 to $100$ pixels), and random dropping of $1/6$ of all the 3D minutiae based feature points.
Appendix B Supplement to the Feature Extraction Experiments
We present minutiae extraction accuracy at 8-pixel resolution in Table 4, and the results demonstrate the same conclusions as those observed at 16-pixel resolution.
Database Method Precision Recall F1-score
CFPose
VeriFinger
55.9% 69.4% 61.4%
Tan and Kumar [tan2020towards]
75.9% 73.3% 74.1%
Proposed
76.9% 70.8% 73.2%